OpenStudy (anonymous):

You want to buy your friend some $50 i-tunes gift cards. You have a $5 off coupon for the total purchase. Write a function to represent the cost of the gifts. What would be the cost of 4 gift cards?

OpenStudy (jdoe0001):

\(\begin{array}{llll} \textit{gift cards}&\textit{cards total cost}&\textit{discount}\\ \hline\\ 2&50\times 2&-5\\ 5&50\times 5&-5\\ 12&50\times 12&-5\\ 20&50\times 20&-5\\ 100&50\times 100&-5\\ \end{array}\) what do you think?

OpenStudy (anonymous):

50 x 4 - 5?

OpenStudy (anonymous):

the function should be represented as \(f(x)=something\) where \(x\) is the number of gift cards purchased. in your case it will be \(f(x)=50x-5\)

OpenStudy (anonymous):

your answer above it the answer to the second part. the cost of 4 gift cards will be \[f(4)=50\times 4-5\]

OpenStudy (anonymous):

Thank you so much!